OpenStudy (anonymous):

\[\sqrt{8x-9}=1\] solve

OpenStudy (anonymous):

8x-9=1 as sqroot1=1 8x-9=1 8x=10 x=(4/5) i think

OpenStudy (anonymous):

if \[\sqrt{8x-9} = 1\] \[\sqrt{8x-9}^{2} = 1^{2}\] \[8x-9 = 1\] \[8x=10\] \[x = 5/4\]

OpenStudy (anonymous):

sorry yes mixed up my fraction spencer is right

OpenStudy (anonymous):

Thanks guys!